OpenStudy (anonymous):

Given a^2=b^2+1, prove that loga+b X=-loga-b X

OpenStudy (anonymous):

given a^2=b^2+1 or a^2-b^2=1 or (a+b)(a-b) =1 taking log of both sides (same base) log (a+b)(a-b) = log 1 or log(a+b) +log(a-b) =0 [as log 1=0 ] or log(a+b) = - log(a-b)

OpenStudy (anonymous):

Thank you ,I'm grateful^^